# BARE: Combining Base and Instruction-Tuned Language Models for Better Synthetic Data Generation

### v2

@@ A.1.1 Fine-Tuning Task Hyperparameters
A.1.2 Classification Task Setup
The generated data is used to train a BERT-based classifier [Devlin et al., 2018] for 2 epochs on Enron and 9 epochs on Newsgroups. The trained models are evaluated on a static test set with $n=500$ examples for each domain.
A.2 Core Experiment Results - All Domains
@@
Table 4: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on on a randomly selected static $n=500$ subset of Newsgroups. A BERT model with a classification head was trained for $9$ epochs on the generated data $(n=500)$.
@@
A.3 Independent Sampling Temperature Ablations - HotpotQA, PubMedQA, and LCB TOP
@@
A.4 BARE First Stage Ablations - GSM8K
This appendix contains diversity, IR, and downstream performance results for our ablation replacing the first stage of BARE with an instruct-tuned model, specifically Llama-3.1-70B-Instruct. We refine using Llama-3.1-70B-Instruct and GPT-4o, and investigate the change in downstream performance compared to standard BARE (using Llama-3.1-70B-Base in the first stage). Note that dataset diversity is unchanged compared to direct generation from Llama-3.1-70B-Instruct, that IR improves after refinement, and that downstream performance is consistently worse than standard BARE.
@@ A.4 BARE First Stage Ablations - GSM8K
Appendix B GSM8K Prompt Examples
@@
Here are some examples: {examples} Now it’s your turn. Generate {num} different problems following this format. Your question should be different in content from the examples. Make sure to only provide only the question and answer. Start each example with the question. Delimit the end of an example with the phrase ”END OF EXAMPLE” (all caps) on a new line.

### v3

@@ B.1.1 Fine-Tuning Task Hyperparameters
B.1.2 Classification Task Setup
The generated data is used to train a BERT-based classifier Devlin et al. [2018] for 2 epochs on Enron and 9 epochs on Newsgroups. The trained models are evaluated on a static test set with $n=500$ examples for each domain.
B.1.3 Compute
We use 1-4 A100s for serving open-source models for data generation tasks (8B and 70B models) and up to a single node of 8xH100s for training purposes. LoRA fine-tuning takes on the order of minutes and data generation for 10,000 examples for an 70B model using 90 cores takes approximately an hour.
B.2 Core Experiment Results - All Domains
@@
Table 4: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a randomly selected static $n=500$ subset of Newsgroups. A BERT model with a classification head was trained for $9$ epochs on the generated data $(n=500)$.
@@
B.3 Independent Sampling Temperature Ablations - HotpotQA, PubMedQA, and LCB TOP
@@
B.4 BARE First Stage Ablations - GSM8K
This appendix contains diversity, IR, and downstream performance results for our ablation replacing the first stage of BARE with an instruction-tuned model, specifically Llama-3.1-70B-Instruct. We refine using Llama-3.1-70B-Instruct and GPT-4o, and investigate the change in downstream performance compared to standard BARE (using Llama-3.1-70B-Base in the first stage). Note that dataset diversity is unchanged compared to direct generation from Llama-3.1-70B-Instruct, that IR improves after refinement, and that downstream performance is consistently worse than standard BARE.
@@ B.4 BARE First Stage Ablations - GSM8K
B.5 BARE Training Set Scale Ablations
This appendix contains downstream performance results for our ablation where we scale up the number of BARE generated data points from $1,000$ to $10,000$ in an effort to investigate whether the trends hold at larger scales. We focused our investigation of this particular ablation on the Llama-3.1-8B model family, and across both the LCB TOP and PubMedQA domains we find that the BARE trends consistently hold at scale.
Table 13: Downstream accuracy results on two evaluation sets (LCB and PubMedQA) using 10,000 BARE-generated data points and the Llama-3.1-8B model family. Models were fine-tuned for 2 epochs rather than the 4 used for 1,000 points due to compute considerations.
Evaluation Set Base Instruct BARE
LCB TOP 0.192 0.181 0.233
PubMedQA 0.512 0.645 0.658
Appendix C Additional Discussion
In this appendix we further discuss our evaluation results and propose hypotheses for certain trends or observations.
Why is BARE generation sometimes worse than instruction-tuned generation?
As mentioned in the main paper, the choice of refiner plays an important role in the BARE pipeline. A model that is proficient at generating data may not be proficient at improving data. Mitigation of this issue can include tailoring prompts for specific models, but we chose to keep our prompts consistent and general to focus on the core idea of leveraging base mode generations.
Why is instruction-tuned generation sometimes worse than base generation?
A key insight to our work is the diversity-quality tradeoff between base models and instruction-tuned models. In some cases, the additional diversity from base models outweighs the lower quality.
Why do smaller models sometimes produce better data than larger models?
The diversity-quality tradeoff between smaller and larger models is not well understood. While it is generally accepted that larger models are of higher quality, it is unclear whether they are similar in terms of diversity. In some cases, the smaller model exhibits higher diversity in its generations, allowing the resulting datasets to be of more utility despite the lower quality.
Appendix D GSM8K Prompt Examples
@@
Here are some examples: {examples} Now it’s your turn. Generate {num} different problems following this format. Your question should be different in content from the examples. Make sure to only provide only the question and answer. Start each example with the question. Delimit the end of an example with the phrase "END OF EXAMPLE" (all caps) on a new line.